OpenStudy (anonymous):

Solve the following system of equations algebraically. Verify your solution either graphically or by using matrices 8x-2y=5 -12x+3y=7

OpenStudy (zpupster):

you have to eliminate an x or y . do that by multiplying one of the equations or both equations by a number that will accomplish that. here you would multiply top equation by 3 and bottom equation by 2. 8x-2y=5 -12x+3y=7 this will give: 24x-6y=15 -24x +6y =14 both x and y cancel there is no solution. this is one way another way is to solve for a variable and then plug your answer back in from the graph you can see no intersection points the line equations are parallel no solutions

OpenStudy (anonymous):

So there is no solution for the equation, that is the answer?

OpenStudy (anonymous):

yes there is no solution for the equation